# Degree two Gopakumar–Vafa invariants of local curves

### v2

@@ 1.1. Motivation and results
In this paper we study another curve counting theory on a CY $3$-fold, that is Gopakumar–Vafa (GV) theory. GV theory was originally introduced in physics [GV98], as a way to count curves in a CY $3$-fold. Recently, Maulik–Toda [MT18] proposed a mathematically rigorous definition of GV invariants, building on the previous works by Hosono–Saito–Takahashi [HST01] and Kiem–Li [KL16]. See the next subsection for the review of Maulik–Toda’s work. GV theory is also expected to be equivalent to GW theory (the GV/GW correspondence conjecture):
@@ 1.2. GV invariants of a local curve and strategy of the proofs
where $\phi_{M}$ on the left hand side denotes Joyce’s perverse sheaf on $M_{X}(\beta,1)$. One of the main difficulties in this approach to GV theory is the complicated gluing procedure of the construction of $\phi_{M}$: it is defined as a gluing of locally defined perverse sheaves of vanishing cycles.
@@ 1.3. Relation to previous work
Shen–Yun [SY22] investigated a compact analogue of the P=W phenomenon for Lagrangian fibrations. As an application, they established the GV/PT correspondence for a local K3 surface.
@@ Acknowledgements.
We would like to thank Jim Bryan for introducing this problem to us. We would also like to thank Tasuki Kinjo, Davesh Maulik and Yukinobu Toda for related discussions.
B.D is supported by a Royal Society university research fellowship.
@@ 2.1. Twisted Higgs bundles
Let $N$ be a rank two vector bundle on $C$ with $\det(N)\cong\omega_{C}$. Choose an exact sequence
@@
We need to prove that the inclusion $i\colon\mathcal{M}_{N}(2,d)\subset\mathcal{M}_{L}(2,d)$ preserves $\mu$-stability. Suppose for a contradiction that there exists $(E,\phi)\in M_{N}^{st}(2,d)$ such that $i(E,\phi)$ is not $\mu$-stable. Then by the proof of [KK23, Lemma 3.4], the maximal destabilizing subsheaf $F\subset E$ satisfies $\operatorname{Hom}(F,E/F\otimes L^{-1}\otimes\omega_{C})\neq 0$. Under the current assumption of $\operatorname{rk}(E)=2$, $F$ and $E/F$ are line bundles with $\deg(F)>\deg(E/F)$. Moreover, we have $\deg(L^{-1}\otimes\omega_{C})\leq 2g(C)-2-(2(g)-1)<0$. This leads to a contradiction. ∎
@@ 2.2. Rigidity of vector bundles
Fix a line bundle $L$ on $C$ with $\deg(L)\geq 2g(C)-1$. Then a general $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $2$-rigid.
@@
By the arguments in [BP06, Section 3], $2$-rigid bundles are in the complement of $\Delta(d)\subset\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ for $d\geq 0$, where
$\Delta(d)\coloneqq\left\{[N]\colon\begin{aligned} &\text{there exists a saturated subbundle }S\subset N\text{ with }\\ &\deg(S)=d,\operatorname{H}^{0}(S^{\otimes 2})\neq 0\end{aligned}\right\}.$
The proof of [BP06, Lemma 3.1] shows that
$\dim\Delta(d)<3g(C)-3<3g(C)-1\leq\dim\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$
@@
Note that as $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is not necessarily slope semistable, $\Delta(d)$ may not be empty for $d\geq g(C)$. Nevertheless, the same argument shows that $\Delta(d)\subset\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is a proper closed subset for any $d\geq 0$, and that $\Delta(d)=\emptyset$ for $d\gg 0$. ∎
@@ 2.3. Hitchin fibrations and Gopakumar–Vafa invariants
Let $s\colon f^{-1}(\mathbb{R}_{>0})\hookrightarrow B_{L}$ and $z\colon f^{-1}(0)\hookrightarrow B_{L}$ be the inclusions. Then by the definition of the nearby cycles functor, and one more application of Lemma 2.10 we have
$\operatorname{H}(B_{L},i_{*}i^{*}\psi_{g}P)=$ $\operatorname{H}(B_{L},i_{*}i^{*}z_{*}z^{*}s_{*}s^{*}P)$
@@
    Computation of $n_{g}(U_{\varepsilon})$: this part can be done by determining the fibre-wise contributions. Since $U_{\varepsilon}$ avoids the origin, the spectral curve $C_{a}$ is reduced at every point of $a\in U_{\varepsilon}$. Hence, the fibrewise contributions are computed by certain HOMFLY polynomials [MS13, Mau16, MY14, OS12]. We further need to calculate the stratification on $U_{\varepsilon}$ induced by the types of singularities of the spectral curves. This last step requires a careful analysis of the geometry of certain degeneracy loci. This analysis becomes increasingly complicated as we allow the genus to increase, and for this reason in this paper we only give the full result in the case $g(C)=2$.
@@ 3.1. Stratification of the Hitchin base
$S_{\lambda}\coloneqq\{\lambda_{1}x_{1}+\cdots\lambda_{m}x_{m}:x_{i}\in C,x_{i}\neq x_{j}\text{ for }i\neq j\}.$
@@
$\overline{S}_{\lambda}=\{\lambda_{1}x_{1}+\cdots\lambda_{m}x_{m}:x_{i}\in C\}=\coprod_{\lambda\vdash\mu}S_{\mu}.$
@@ 3.1.1. $\lambda=(6)$
$C\cong\Theta\subset\operatorname{Pic}^{1}(C)\xhookrightarrow{(-)^{\otimes 6}}\operatorname{Pic}^{6}(C),$
@@ 3.1.4. $\lambda=(4,1,1)$
as follows. The divisor $H_{\infty}$ is linearly equivalent to the divisor $D_{p}$ defined by the vanishing of the section $s$ at $p$, for some fixed $p\in C$. If a section $s=\psi(q)$ in $Z$ vanishes at $p$, either $p=x$, in which case the intersection multiplicity is $4$, or $p$ is one of the other two zeros of $\Psi(q)$. Counting these points corresponds to counting solutions to $L^{\otimes 2}(-p)\cong\mathcal{O}_{C}(4x+y)$. But now this is the same theta divisor calculation as we have seen before, and we find that there are $24^{2}*1^{2}*$ 2 such points. So $Z\cap H_{\infty}=32+4=36$.
@@ 3.1.6. $\lambda=(3,2,1)$
Let $Z\subset C\times C\times C$ be the subvariety defined by $\mathcal{O}_{C}(3+2y+z)\cong L^{\otimes 2}$. We claim that $e(Z)=-196$. We postpone the proof until the next subsection, see Proposition 3.6.
Fixing some $p\in C$, the preimage $D$ of the divisor $H_{\infty}$ is equivalent to the divisor $3\{p\}\times C\times C+2C\times\{p\}\times C+C\times C\times\{p\}$. So $D$ consists of $132=3*8+2*18+72$ points. These numbers come from the same calculation of intersections of theta divisors as before, so for example $72$ appears via $72=2*2^{2}*3^{2}$. $Z$ also contains some copies of the deeper strata: it contains one copy of $S_{(4,2)}$ (corresponding to $z=x$), one copy of $S_{(5,1)}$ given by setting $x=y$, and two copies of $S_{(3,3)}$ (corresponding to $y=z$). So putting everything together we find
@@ 3.1.7. $\lambda=(3,1^{3})$
Next we remove deeper strata: one copy of $S_{(3,2,1)}$, one copy of $S_{(4,1,1)}$, two copies of $S_{(3,3)}$, one copy of $S_{(4,2)}$, and one copy of $S_{(5,1)}$. Putting everything together we find
@@ 3.1.9. $\lambda=(2,2,1,1)$
Consider a rational map $\phi\colon C\times C\dashrightarrow\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$ which sends generic $(x,y)\in C\times C$ to the one-dimensional subspace $\operatorname{H}^{0}(L^{\otimes 2}(-2x-2y))\subset\operatorname{H}^{0}(L^{\otimes 2})$. As discussed above, the indeterminancy locus of this rational map is $32$ points in $C\times C$. In the following, we calculate the line bundle inducing the rational map $\phi$. Consider the following diagram:
@@
In all, we must remove three copies of $S_{(4,2)}$, two copies of $S_{(3,3)}$, two copies of $S_{(3,2,1)}$, six copies of $S_{(2,2,2)}$, one copy of $S_{(5,1)}$, and one copy of $S_{(4,1,1)}$. The number six deserves some explanation. Let $(a,b,c)$ be a point of $\operatorname{Sym}^{6}(C)$, lying in the image of $S_{(2,2,2)}$. Then we can lift it to a point of $Z$ in six different ways, the first by setting $x=a,y=b,z=w=c$, and the other five by permuting $a,b,c$ in these three equalities. Finally, note that after removing $T$, the union of all of these contributions from deeper strata, the cover $Z\setminus T\rightarrow S_{(2,2,1,1)}$ is $2:1$. Putting this all together, we find
@@ 3.1.10. $\lambda=(2,1^{4})$
$2S_{(2,2,1,1)}\cup S_{(3,1,1,1)}\cup 3S_{(2,2,2)}+2S_{(3,2,1)}+S_{(4,1,1)}+2S_{(3,3)}+2S_{(4,2)}+S_{(5,1)}$
The above coefficients are given as follows: a point in, say $S_{(3,2,1)}$, is represented by a tuple $(x_{0},\ldots,x_{5})\in\operatorname{Sym}^{6}(C)$ with $x_{0}=x_{1}=x_{2}$ and $x_{3}=x_{4}$. In order to lift to $Z$ we need to pick one out of $x_{0}$ or $x_{3}$ to be the value of $x$. In general the coefficient of $S_{\pi}$ is given by counting the number of rows in $\pi$ of length at least $2$.
@@ 3.2. A degeneracy locus calculation
By abuse of notation, we denote the pushforward via $\mu$ of the above classes by the same symbols.
We first compute the Chern character of $\mathcal{F}$ by using the Grothendieck–Riemann–Roch formula:
@@
    $\operatorname{ch}_{4}(\mathcal{F})=66\operatorname{pt}$.
@@
The first two equalities follow by observing that $\Delta_{1}=p_{01}^{*}\Delta$, where $p_{01}$ is the projection to the first two factors, and $\Delta\subset C\times C$ denotes the diagonal. Then the third one follows as $\delta=\Delta_{1}.\Delta_{23}$.
@@
$c_{2}(\mathcal{F})=\frac{1}{2}\operatorname{ch}_{1}(\mathcal{F})^{2}-\operatorname{ch}_{2}(\mathcal{F});\quad c_{3}(\mathcal{F})=\frac{1}{6}\operatorname{ch}_{1}(\mathcal{F})^{3}-\operatorname{ch}_{1}(\mathcal{F})\operatorname{ch}_{2}(\mathcal{F})+\frac{1}{6}\operatorname{ch}_{1}(\mathcal{F})^{3}.$

### v3

@@ 1.1. Motivation and results
In this paper we study another curve counting theory on a CY $3$-fold, that is Gopakumar–Vafa (GV) theory. GV theory was originally introduced in physics [GV98], as a way to count curves in CY $3$-folds. Recently, Maulik–Toda [MT18] proposed a mathematically rigorous definition of GV invariants, building on the previous works by Hosono–Saito–Takahashi [HST01] and Kiem–Li [KL16]. See the next subsection for the review of Maulik–Toda’s work. GV theory is also expected to be equivalent to GW theory (the GV/GW correspondence conjecture):
@@ 1.2. GV invariants of a local curve and strategy of the proofs
where $\phi_{M}$ on the left hand side denotes Joyce’s perverse sheaf on $M_{X}(\beta,1)$. Here, and throughout, if $\mathcal{P}$ is a constructible complex of sheaves on a space $T$, we define $\chi(\mathcal{P})\coloneqq\sum_{i\in\mathbb{Z}}(-1)^{i}\dim(\operatorname{H}^{i}(T,\mathcal{P}))$.One of the main difficulties in this approach to GV theory is the complicated gluing procedure of the construction of $\phi_{M}$: it is defined as a gluing of locally defined perverse sheaves of vanishing cycles.
@@ 1.3. Relation to previous work
Shen–Yin [SY22] investigated a compact analogue of the P=W phenomenon for Lagrangian fibrations. As an application, they established the GV/PT correspondence for a local K3 surface.
@@ Acknowledgements.
We would like to thank Jim Bryan for introducing this problem to us. We would also like to thank Tasuki Kinjo, Davesh Maulik and Yukinobu Toda for related discussions. We also thank the anonymous referee for many helpful suggestions and corrections.
B.D is supported by a Royal Society university research fellowship
[URF/R/221040].
@@ 2.1. Twisted Higgs bundles
[KM24, Theorem 5.6, Proposition 5.7] Let $N$ be a rank two vector bundle on $C$ with $\det(N)\cong\omega_{C}$. Choose an exact sequence
@@
We need to prove that the inclusion $i\colon\mathcal{M}_{N}(2,d)\subset\mathcal{M}_{L}(2,d)$ preserves $\mu$-stability. Suppose for a contradiction that there exists $(E,\phi)\in M_{N}^{st}(2,d)$ such that $i(E,\phi)$ is not $\mu$-stable. Then by the proof of [KK23, Lemma 3.4], the maximal destabilizing subsheaf $F\subset E$ satisfies $\operatorname{Hom}(F,E/F\otimes L^{-1}\otimes\omega_{C})\neq 0$. Under the current assumption of $\operatorname{rk}(E)=2$, $F$ and $E/F$ are line bundles with $\deg(F)>\deg(E/F)$. Moreover, we have $\deg(L^{-1}\otimes\omega_{C})\leq 2g(C)-2-(2g(C)-1)<0$. This leads to a contradiction. ∎
@@ 2.2. Rigidity of vector bundles
Fix a line bundle $L$ on $C$ with degree $d\geq 2g(C)-1$. Assume that $L$ is a generic element in $\operatorname{Pic}^{d}(C)$. Then a general $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $2$-rigid.
@@
By the arguments in [BP06, Section 3], stable bundles in the complement of $\Delta(d)\subset\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ for $d\geq 0$, where
$\Delta(d)\coloneqq\left\{[N]\colon\begin{aligned} &\text{there exists a saturated subbundle }S\subset N\text{ with }\\ &\deg(S)=d,\operatorname{H}^{0}(S^{\otimes 2})\neq 0\end{aligned}\right\},$
are $2$-rigid. The proof of [BP06, Lemma 3.1] shows that
$\dim\Delta(d)<3g(C)-3<3g(C)-2\leq\dim\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$
@@
It remains to show that a generic $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is slope semistable, which follows from Lemma 2.7 below. ∎
An earlier version of this paper omitted the check that generic $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is slope semistable; we thank the anonymous referee for pointing this out, as well as suggesting the proof of the following lemma.
Lemma 2.7.
Let $L$ be a line bundle of degree $d\geq g(C)-1$. Assume that $L$ is a generic element in $\operatorname{Pic}^{d}(C)$. Then a general $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $\mu$-semistable.
Proof.
Since $\mu$-semistability is an open condition, it is enough to find a line bundle $L\in\operatorname{Pic}^{d}(C)$ and a point $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ such that $N$ is $\mu$-semistable.
We may find a line bundle $L_{0}\in\operatorname{Pic}^{d}(C)$ and line bundles $M_{0},M_{1}$ of degree $g(C)-1$ such that there is a surjection $M_{0}\oplus M_{1}\to L_{0}$. We may then take a line bundle $L^{\prime}\in\operatorname{Pic}^{0}(C)$ such that $M_{0}\otimes M_{1}\otimes(L^{\prime})^{\otimes 2}\cong\omega_{C}$. Now, set $N\coloneqq(M_{0}\oplus M_{1})\otimes L^{\prime}$ and $L\coloneqq L_{0}\otimes L^{\prime}$. Since $M_{0}$ and $M_{1}$ have the same degree, $N$ is $\mu$-semistable. Furthermore, we have $L\in\operatorname{Pic}^{d}(C)$ and there exists a short exact sequence
$0\to L^{-1}\otimes\omega_{C}\to N\to L\to 0$
as required. ∎
@@ 2.3. Hitchin fibrations and Gopakumar–Vafa invariants
Let $s\colon f^{-1}(\mathbb{R}_{>0})\hookrightarrow B_{L}$ and $z\colon f^{-1}(0)\hookrightarrow B_{L}$ be the inclusions. Recall also that we denote $i\colon 0\hookrightarrow B_{L}$ the inclusion. Then by the definition of the nearby cycles functor, and one more application of Lemma 2.11 we have
$\operatorname{H}(B_{L},i_{*}i^{*}\psi_{f}P)=$ $\operatorname{H}(B_{L},i_{*}i^{*}z_{*}z^{*}s_{*}s^{*}P)$
@@
    Computation of $n_{g}(U_{\varepsilon})$: this part can be done by determining the fibre-wise contributions. Since $U_{\varepsilon}$ avoids the origin, the spectral curve $C_{a}$ is reduced at every point of $a\in U_{\varepsilon}$. Hence, the fibrewise contributions are computed by certain HOMFLY polynomials [MS13, Mau16, MY14, OS12]. We further need to calculate the stratification of $U_{\varepsilon}$ induced by the types of singularities of the spectral curves. This last step requires a careful analysis of the geometry of certain degeneracy loci. This analysis becomes increasingly complicated as we allow the genus to increase, and for this reason in this paper we only give the full result in the case $g(C)=2$.
@@ 3.1. Stratification of the Hitchin base
$S_{\lambda}\coloneqq\{\lambda_{1}x_{1}+\cdots+\lambda_{m}x_{m}:x_{i}\in C,x_{i}\neq x_{j}\text{ for }i\neq j\}.$
@@
$\overline{S}_{\lambda}=\{\lambda_{1}x_{1}+\cdots+\lambda_{m}x_{m}:x_{i}\in C\}=\coprod_{\lambda\vdash\mu}S_{\mu}.$
@@
We have used the fact that if $S^{\prime}\rightarrow S$ is an unramified $n:1$ cover of a variety $S$, then $\chi(S^{\prime})=n\chi(S)$. In what follows we will denote such $S^{\prime}$ by $S^{[n\operatorname{\!\colon\!\!\!}1]}$; although there may be many such covers of $S$, the ambiguity in the notation will be irrelevant as long as we are only interested in the Euler characteristic of $S^{\prime}$.
@@
    As in the strategy explained above, we then remove $D^{\prime}\subset Z^{\prime}$, the pull back of $H_{\infty}$. It will turn out that $D^{\prime}$ is also a projective space bundle over $C^{\times(m-k)}$, one dimension less than $Z^{\prime}$. Finally, we will remove the deeper strata from $Z^{\prime}$ by the same formula as above.
@@ 3.1.1. $\lambda=(6)$
$C\cong\Theta\subset\operatorname{Pic}^{1}(C)\xrightarrow{(-)^{\otimes 6}}\operatorname{Pic}^{6}(C),$
@@ 3.1.4. $\lambda=(4,1,1)$
as follows. The divisor $H_{\infty}$ is linearly equivalent to the divisor $D_{p}$ defined by the vanishing of the section $s$ at $p$, for some fixed $p\in C$. If a section $s=\psi(q)$ in $Z$ vanishes at $p$, either $p=x$, in which case the intersection multiplicity is $4$, since the regular function defining $D_{p}$ vanishes to order 4,or $p$ is one of the other two zeros of $\Psi(q)$. Counting these points corresponds to counting solutions to $L^{\otimes 2}(-p)\cong\mathcal{O}_{C}(4x+y)$. But now this is the same theta divisor calculation as we have seen before, and we find that there are $24^{2}*1^{2}*$ 2 such points. So $Z\cap H_{\infty}=32+4=36$.
@@ 3.1.6. $\lambda=(3,2,1)$
Let $Z\subset C\times C\times C$ be the subvariety defined by $\mathcal{O}_{C}(3x+2y+z)\cong L^{\otimes 2}$. We claim that $e(Z)=-196$. We postpone the proof until the next subsection, see Proposition 3.6.
Fixing some $p\in C$, the preimage $D$ of the divisor $H_{\infty}$ is equivalent to the divisor $3\{p\}\times C\times C+2C\times\{p\}\times C+C\times C\times\{p\}$. So $D$ consists of $132=3*8+2*18+72$ points. These numbers come from the same calculation of intersections of theta divisors as before, so for example $72$ appears via $72=2*2^{2}*3^{2}$. $Z$ also contains some copies of the deeper strata: it contains one copy of $S_{(4,2)}$ (corresponding to $z=x$), one copy of $S_{(5,1)}$ given by setting $x=y$, and a variety $S^{[2\operatorname{\!\colon\!\!\!}1]}_{(3,3)}$ (corresponding to $y=z$). So putting everything together we find
@@ 3.1.7. $\lambda=(3,1^{3})$
Next we remove deeper strata: one copy of $S_{(3,2,1)}$, one copy of $S_{(4,1,1)}$, a variety $S^{[2\operatorname{\!\colon\!\!\!}1]}_{(3,3)}$, one copy of $S_{(4,2)}$, and one copy of $S_{(5,1)}$. Putting everything together we find
@@ 3.1.9. $\lambda=(2,2,1,1)$
Consider the rational map $\phi\colon C\times C\dashrightarrow\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})$ which sends generic $(x,y)\in C\times C$ to the one-dimensional subspace $\operatorname{H}^{0}(L^{\otimes 2}(-2x-2y))\subset\operatorname{H}^{0}(L^{\otimes 2})$. As discussed above, the indeterminancy locus of this rational map is $32$ points in $C\times C$. In the following, we calculate the line bundle inducing the rational map $\phi$. Consider the following diagram:
@@
In all, we must remove three copies of $S_{(4,2)}$, a double cover $S^{[2\operatorname{\!\colon\!\!\!}1]}_{(3,3)}$, two copies of $S_{(3,2,1)}$, a cover $S^{[6\operatorname{\!\colon\!\!\!}1]}_{(2,2,2)}$, one copy of $S_{(5,1)}$, and one copy of $S_{(4,1,1)}$. The number six deserves some explanation. Let $(a,b,c)$ be a point of $\operatorname{Sym}^{6}(C)$, lying in the image of $S_{(2,2,2)}$. Then we can lift it to a point of $Z$ in six different ways, the first by setting $x=a,y=b,z=w=c$, and the other five by permuting $a,b,c$ in these three equalities. Finally, note that after removing $T$, the union of all of these contributions from deeper strata, the cover $Z\setminus T\rightarrow S_{(2,2,1,1)}$ is $2:1$. Putting this all together, we find
@@ 3.1.10. $\lambda=(2,1^{4})$
$S^{[2\operatorname{\!\colon\!\!\!}1]}_{(2,2,1,1)}+S_{(3,1,1,1)}+S^{[3\operatorname{\!\colon\!\!\!}1]}_{(2,2,2)}+2S_{(3,2,1)}+S_{(4,1,1)}+S^{[2\operatorname{\!\colon\!\!\!}1]}_{(3,3)}+2S_{(4,2)}+S_{(5,1)}$
The degree of a cover $mS^{[n\operatorname{\!\colon\!\!\!}1]}_{\pi}\rightarrow S_{\pi}$ is $mn$. The above degrees are given as follows: a point in, say $S_{(3,2,1)}$, is represented by a tuple $(x_{0},\ldots,x_{5})\in\operatorname{Sym}^{6}(C)$ with $x_{0}=x_{1}=x_{2}$ and $x_{3}=x_{4}$. In order to lift to $Z$ we need to pick one out of $x_{0}$ or $x_{3}$ to be the value of $x$. In general the degree of the cover over $S_{\pi}$ is given by counting the number of rows in $\pi$ of length at least $2$.
@@ 3.2. A degeneracy locus calculation
By abuse of notation, we denote the pushforward via $\mu$ of the above classes by the same symbols. We first compute the Chern character of $\mathcal{F}$ by using the Grothendieck–Riemann–Roch formula:
@@
    $\operatorname{ch}_{3}(\mathcal{F})=66\operatorname{pt}$.
@@
The first two equalities follow by observing that $\Delta_{1}=p_{01}^{*}\Delta$, where $p_{01}$ is the projection to the first two factors, and $\Delta\subset C\times C$ denotes the diagonal. Then the third one follows, as $\delta=\Delta_{1}.\Delta_{23}$.
@@
$c_{2}(\mathcal{F})=\frac{1}{2}\operatorname{ch}_{1}(\mathcal{F})^{2}-\operatorname{ch}_{2}(\mathcal{F});\quad c_{3}(\mathcal{F})=\frac{1}{6}\operatorname{ch}_{1}(\mathcal{F})^{3}-\operatorname{ch}_{1}(\mathcal{F})\operatorname{ch}_{2}(\mathcal{F})+2\operatorname{ch}_{3}(\mathcal{F}).$